QUESTION 9 Tims Petrochemicals produces a variety of industrial chemicals. Bob runs a fish farm downstream from Tims Petrochemicals plant. Currently Tim dumps waste into the river. This pollution causes mutations that result in fish becoming deformed. Bob is unable to sell deformed fish. Bob wishes to come to an agreement with Tim to stop him from polluting the river. Pollution 800 600 1400 Pollution Controls 400 1,102 Gains to Tim (S) Total (S) If negotiation costs are $137 what is the most that Bob could offer Tim to introduce pollution controls and still be better off by $1? Answer to the nearest whole number in dollars (with no decimal places, S sign, spaces or commas).
Answer #1

Total social costs with pollution = $1400

Total social costs with pollution controls = $400 + $ 1102 = $1502

Loss to Tim = 800 - 400 = $400

Gain to Tim = 1102 -600 = $502

Negotiation cost = 132

Net gain to Bob after subtracting negotiation costs = 502 - 132 = $370

Bob would offer the amount $370 - $1 = $369 so as to still make a gain of $1.

Answer is $369.